40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6736     Entry time: Fri Jun 1 02:13:00 2012
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Attempt at Ygreen beat - failed 

[Yuta, Jenne]

We tried to find the Ygreen beat note, with no success yet.  We calculate from Bryan's formula that the Yend laser should be ~34.68C.  But Katrin has an elog saying that she was looking around 19C.  I don't know why the discrepancy, but maybe this is part of our problem?  Kiwamu elog-responded that the epics output had to be high (~9V) when the temp was 19C.  So maybe we need a smaller offset setting in the slow servo with the 34C temperature?

We set the "T+" on the Ygreen laser controller to 34.68C using the dial, and then tried a few large steps with the offset in the Ygreen slow servo.  The idea was to see if we could swing past the beat, so we would know vaguely where it was.  But we never saw a resonance on the spectrum analyzer, even with a "hold max" trace. 

We confirmed that there is signal going to the SLOW input of the laser controller's front panel.  Yuta watched a voltmeter while I changed the epics value, and we successfully changed the signal.  However, after plugging the SLOW cable back in, we noticed that no matter what we set the epics value to, we don't see any temperature change reported on the front panel display.  There is something in the manual( according to Katrin) that the "LT" display is not accurate when a cable is plugged in.  But none of the display values changed.  I think there is a measured temp output on the back that Bryan mentioned that we could use to see if something is really changing inside.

Anyhow, no beatnote found yet tonight.  We confirmed before starting that the alignment onto the beat PD was good, so that's not the problem.

ELOG V3.1.3-